# Dimension of Inhomogeneous Sub-Self-Similar Sets

### v1

@@ Abstract.
In this paper, we introduce the concept of Inhomogeneous sub-self-similar (ISSS) sets, building upon the foundations laid by Falconer (Trans. Amer. Math. Soc. 347 (1995) 3121–3129) in the study of sub-self-similar sets and drawing inspiration from Barnsley’s work on inhomogeneous self-similar sets (Proc. Roy. Soc. London Ser. A 399 (1985), no. 1817, 24). We explore a range of examples of ISSS sets and elucidate a method to construct ISSS sets. We also investigate the upper and lower box dimensions of ISSS sets and discuss the continuity of the Hausdorff dimension.
Key words and phrases:
@@ 1. Introduction
A wide variety of fractals exhibit self-similarity, i.e., they consist of infinitely smaller replicas that resemble the whole. Hutchinson [20] provided a comprehensive framework for self-similar sets and constructed them mathematically using Iterated Function Systems (IFS). The construction of self-similar sets through IFS has been extensively studied and popularized by Barnsley and his collaborators [3, 5, 4]. Barnsley [3] introduced the concept of fractal interpolation functions (FIF), and many researchers have since worked on FIFs and their applications. For more details, interested readers may see [8, 7, 23, 26, 32, 35, 37, 36, 21]. Barnsley and Demko [4] extended the IFS framework by introducing inhomogeneous IFS, which includes traditional IFS along with a compact set called a condensation set. Inhomogeneous IFS generalizes the concept of IFS in such a way that, by considering the condensation set as the empty set, it reduces to the classical self-similar case. Furthermore, Mauldin and Williams [30] introduced the graph-directed IFS (GD-IFS), which generalizes traditional IFS. In this regard, Boore and Falconer [6] provide an example of a GD fractal that cannot be generated by a traditional IFS. Recently, Dubey and Verma further generalized GD-IFS by introducing the concept of inhomogeneous GD-IFS (see [10, 11, 33] for more details). In [14], Falconer introduced the concept of sub-self-similar (SSS) sets by replacing equality in the self-similar equation with set inclusion. Computing the exact dimension of a fractal set remains a challenging task in fractal geometry. Various methods exist for estimating the dimensions of fractals. Chandra and Abbas [9] computed the fractal dimension of mixed Riemann–Liouville integrals using the $\delta$-covering method. Falconer [12] discussed fractal dimensions using the Fourier transform method. Priyadarshi [34] estimated the lower bound on the Hausdorff dimension of sets of complex continued fractions using operator-theoretic techniques. Chandra and Abbas [8] studied the fractal dimension of fractal functions using function space techniques. Falconer [12] also explored fractal dimensions using potential-theoretic techniques and the mass distribution principle. Hochman [19] investigated the Hausdorff dimension of self-similar sets and measures on the real line using entropy techniques and provided an inverse theorem for the entropy of convolution. Shmerkin [40] extended this work by studying the $L^{q}$-dimension of dynamically driven self-similar measures. Following Falconer’s work [14] and using the concept of inhomogeneous self-similar sets, we introduce and construct inhomogeneous sub-self-similar sets (ISSS). Motivated by Fraser’s work [15], we estimate the fractal dimension of these sets using the $\delta$-covering method.
@@
The outline of the paper is as follows: Section 2 provides the preliminaries. In Section 3, we present examples of ISSS sets. Section 4 discusses one of the methods for constructing ISSS sets, while Sections 5 and 6 are devoted to estimating the fractal dimension of these sets and discussing the continuity of the Hausdorff dimension for ISSS sets. In Section 7, we study the self-similar structure on the product of inhomogeneous IFS. Finally, Section 8 concludes the paper with remarks and outlines potential future directions for research.
@@ 2.1. Code Space
Let $I^{\infty}$ denote the set of all infinite sequences in $\{1,\ldots,N\}$, that is,
@@
An element of $I^{\infty}$ is written as $\omega:=\omega_{1}\omega_{2}\dots\omega_{k}\omega_{k+1}\dots$, and the $k$-th term of the sequence is denoted by $\omega_{k}$.
@@
$S^{n}:=\left\{\omega_{|_{n}}=\omega_{1}\dots\omega_{n}\,\middle|\,\omega=\omega_{1}\omega_{2}\dots\omega_{k}\dots\in S\right\},\quad\text{and}\quad S^{*}:=\bigcup_{n}S^{n}.$
@@
A compact subset $K\subseteq I^{\infty}$ is said to be closed under left shift if, for every $\omega=\omega_{1}\omega_{2}\dots\in K$, the sequence $\omega^{\prime}:=\omega_{2}\omega_{3}\dots$ also belongs to $K$. Falconer characterizes sub-self-similar sets and establishes a natural correspondence between such sets and subsets of the code space. For more details, see [14].
@@ 3. Examples and Construction of Inhomogeneous Sub-Self-Similar Sets
Unlike inhomogeneous self-similar sets, ISSS sets are not unique. For example, consider the IFS $\mathcal{I}=\{f_{1}(x)=\frac{x}{2},f_{2}(x)=\frac{x}{2}+\frac{1}{2}\}$ with condensation set $C$. Both the set $\{0,1\}$ and the set of all dyadic numbers in $[0,1]$ are ISSS sets corresponding to $\mathcal{I}$ and $C$.
@@
In [31], McClure and Vallin provide examples of non-sub-self-similar sets, such as $E=\{0,1,\frac{1}{2},\frac{1}{3},\ldots\}$, which is not a sub-self-similar set. According to the previous remark, there exists a condensation set $C$ such that $E$ becomes an ISSS set. Furthermore, for any compact set $E$, the minimal condensation set $C=\overline{E\setminus\bigcup_{i=1}^{N}f_{i}(E)}$ ensures that $E$ is an ISSS set.
@@ 4. Construction of Inhomogeneous Sub-Self-Similar Sets
$O_{S}=\bigcup_{{\omega}\in S^{*}}f_{{\omega}}(C).$
@@ 5.1. Hausdorff dimension
Let $E\cup O_{S}$ be an ISSS set with SSS set $E$ and condensation set $C.$ Then $E\cup O_{S}=\max\{s,\dim_{H}C\}$.
@@ 5.2. Upper box dimension
The function $\tau(s)=\lim_{k\to\infty}\bigg{(}\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\bigg{)}^{1/k}$ defined in Theorem (5.1) is strictly decreasing continuous function with $0\leq\tau(s)<\infty$ and $\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\geq\tau(s)^{k}$ for all $k=1,2,\ldots$
@@
By Theorem (5.1) $\overline{\dim}_{B}E\leq s,$ it is sufficient to show that $\overline{\dim}_{B}O_{S}\leq\max\{s,\overline{\dim}_{B}C\}.$ Let $\max\{s,\overline{\dim}_{B}C\}<t,$ then by the definition of the box-dimension, there exists some constant $c_{t}$ such that
@@
The third line comes from Lemma (5.6). The sixth line comes as $X$ is a compact space and $\delta/\rho_{\omega}>1,$ the last line arrived by Lemma (5.4). This completes the proof. ∎
@@ 5.3. Lower box dimension
By Lemma (5.4), for $t>s$, the cardinality of the $\delta$-stopping set $S$ is given by:
@@
Let $t>\max\{s,\underline{\dim}_{B}C\}$. Then, by Lemma (5.8), we have $P_{t}(C)\leq\underline{\dim}_{B}C<t$. Thus, by Lemma (5.9), for all $\epsilon\in(0,1-P_{t}(C))$, there exists $\delta\in(0,\epsilon)$ such that:
@@ 6. Continuity of the Hausdorff dimension of ISSS
Let $I_{k}=\{f_{\omega}:\omega\in S^{k}\}$. For any $x\in F=E\cup O_{S}$, we consider two cases:
@@
For each $k\in\mathbb{N}$, define $\omega_{j}|_{k}^{k}:=\omega_{jk+1}\omega_{jk+2}\ldots\omega_{jk+k}$ for $j=0,1,2,\ldots$. Then $\omega_{j}|_{k}^{k}\in S^{k}$, and we have
$x=\lim_{j\to\infty}f_{\omega_{0}|_{k}^{k}}\circ f_{\omega_{1}|_{k}^{k}}\circ\ldots\circ f_{\omega_{j}|_{k}^{k}}(z),$
@@
Conversely, suppose $x\in\bigcup_{k=1}^{\infty}F_{k}\cup O_{S}$. If $x\in O_{S}$, then the claim holds trivially. Now suppose $x\in F_{k}$ for some $k$. Then there exists a sequence $\omega_{0}|_{k}^{k},\omega_{1}|_{k}^{k},\ldots,\omega_{j}|_{k}^{k}\in S^{k}$ such that
$x=\lim_{j\to\infty}f_{\omega_{0}|_{k}^{k}}\circ f_{\omega_{1}|_{k}^{k}}\circ\ldots\circ f_{\omega_{j}|_{k}^{k}}(z).$
Using the fact that $\Theta(S)=E$ and that the concatenation $\omega=\omega_{0}|_{k}^{k}\omega_{1}|_{k}^{k}\ldots\omega_{j}|_{k}^{k}\in S$, we conclude that $x\in F$.
@@
Let $F=E\cup O_{S}$ be the ISSS set corresponding to the IFS $\{f_{i}\}{i=1}^{N}.$ Then there exists a sequence of IFSs $\{I_{k}\}$ with ISSS sets $\{F_{k}\}$ and under inhomogeneous open set condition (IOSC) $\dim_{H}E=\lim_{k\to\infty}\dim_{H}E_{k}=\lim_{k\to\infty}=\lim_{k\to\infty}s_{k},$ where $s_{k}$ is the Hausdorff dimension of $F_{k}$ defined in Lemma 6.2.
@@
By Theorem 6.1, there exists a sequence of IFSs $\{I_{k}\}$ with ISSS sets $F_{k}$ such that $F=\bigcup_{k=1}^{\infty}F_{k}.$ It is easy to observe that if $\{f_{i}\}_{i=1}^{N}$ satisfies the IOSC, then each $I_{k}$ also satisfies the IOSC. If $s_{k}$ represents the Hausdorff dimension of $E_{k}$, which satisfies

### v2

@@ Abstract.
In this paper, we introduce the concept of Inhomogeneous sub-self-similar (ISSS) sets, building upon the foundations laid by Falconer (Trans. Amer. Math. Soc. 347 (1995) 3121–3129) in the study of sub-self-similar sets and drawing inspiration from Barnsley’s work on inhomogeneous self-similar sets (Proc. Roy. Soc. London Ser. A 399 (1985), no. 1817, 24). We explore a range of ISSS sets and present a method for constructing them. We also investigate the upper and lower box dimensions of ISSS sets and discuss the continuity of the Hausdorff dimension.
@@ 1. Introduction
A wide variety of fractals exhibit self-similarity, i.e., they consist of infinitely smaller replicas that resemble the whole. Hutchinson [20] provided a comprehensive framework for self-similar sets and constructed them mathematically using Iterated Function Systems (IFS). The construction of self-similar sets through IFS has been extensively studied and popularized by Barnsley and his collaborators [3, 5, 4]. Barnsley [3] introduced the concept of fractal interpolation functions (FIF), and many researchers have since worked on FIFs and their applications. For more details, interested readers may see [8, 7, 23, 26, 32, 35, 37, 36, 21]. Barnsley and Demko [4] extended the IFS framework by introducing inhomogeneous IFS, which includes traditional IFS along with a compact set called a condensation set. Inhomogeneous IFS generalizes the concept of IFS in such a way that, by considering the condensation set as the empty set, it reduces to the classical self-similar case. Furthermore, Mauldin and Williams [30] introduced the graph-directed IFS (GD-IFS), which generalizes traditional IFS. In this regard, Boore and Falconer [6] provide an example of a GD fractal that cannot be generated by a traditional IFS. Recently, Dubey and Verma further generalized GD-IFS by introducing the concept of inhomogeneous GD-IFS (see [10, 11, 33] for more details). In [14], Falconer introduced the concept of sub-self-similar (SSS) sets by replacing equality in the self-similar equation with set inclusion. Computing the exact dimension of a fractal set remains a challenging task in fractal geometry. Various methods exist for estimating the dimensions of fractals. Chandra and Abbas [9] computed the fractal dimension of mixed Riemann–Liouville integrals using the $\delta$-covering method. Falconer [12] discussed fractal dimensions using the Fourier transform method. Priyadarshi [34] estimated the lower bound on the Hausdorff dimension of sets of complex continued fractions using operator-theoretic techniques. Chandra and Abbas [8] studied the fractal dimension of fractal functions using function space technique. Falconer [12] also explored fractal dimensions using potential-theoretic techniques and the mass distribution principle. Hochman [19] investigated the Hausdorff dimension of self-similar sets and measures on the real line using entropy technique and provided an inverse theorem for the entropy of convolution. Shmerkin [40] extended this work by studying the $L^{q}$-dimension of dynamically driven self-similar measures. Following Falconer’s work [14] and using the concept of inhomogeneous self-similar sets, we introduce and construct inhomogeneous sub-self-similar sets (ISSS). Motivated by Fraser’s work [15], we estimate the fractal dimension of these sets using the $\delta$-covering method.
@@
The outline of the paper is as follows: Section 2 provides the preliminaries. In Section 3, we present examples of ISSS sets. Section 4 discusses a method for constructing ISSS sets, while Sections 5 and 6 are devoted to estimating the fractal dimension of these sets and examining the continuity of the Hausdorff dimension for ISSS sets, respectively. In Section 7, we study the self-similar structure on the product of inhomogeneous IFS. Finally, Section 8 concludes the paper with remarks and outlines for potential future research directions.
@@ 2.1. Code Space
Let $I^{\infty}$ denotes the set of all infinite sequences in $\{1,\ldots,N\}$, that is,
@@
We write an element of $I^{\infty}$ as $\omega:=\omega_{1}\omega_{2}\dots\omega_{k}\omega_{k+1}\dots$, and the $k$-th term of the sequence is denoted by $\omega_{k}$.
@@
$S^{n}:=\left\{\omega_{|_{n}}=\omega_{1}\dots\omega_{n}\,\middle|\,\exists~\omega=\omega_{1}\omega_{2}\dots\omega_{k}\dots\in S\right\},\quad\text{and}\quad S^{*}:=\bigcup_{n}S^{n}.$
@@
A compact subset $K\subseteq I^{\infty}$ is said to be closed under left shift if for every $\omega=\omega_{1}\omega_{2}\dots\in K$, the sequence $\omega^{\prime}:=\omega_{2}\omega_{3}\dots$ also belongs to $K$. Falconer characterizes sub-self-similar sets and establishes a natural correspondence between such sets and subsets of the code space. For more details, see [14].
@@ 3. Examples and Construction of Inhomogeneous Sub-Self-Similar Sets
Unlike inhomogeneous self-similar sets, ISSS sets are not unique. For example, consider the IFS $\mathcal{I}=\{f_{1}(x)=\frac{x}{2},f_{2}(x)=\frac{x}{2}+\frac{1}{2}\}$ with a condensation set $C$. Both the set $\{0,1\}$ and the set of all dyadic numbers in $[0,1]$ are ISSS sets corresponding to $\mathcal{I}$ and $C$. In particular, if we consider $C=\{\frac{1}{2}\},$ then set $\{\frac{1}{2^{n}}:n\geq 1\}\cup\{1+\frac{1}{2^{n}}:n\geq 1\}$ is an ISSS with condensation set $C$ set but not an SSS set.
@@
In [31], McClure and Vallin provide examples of non-sub-self-similar sets, such as $E=\{0,1,\frac{1}{2},\frac{1}{3},\ldots\}$, which is not a sub-self-similar set. As noted previously, there exists a condensation set $C$ such that $E$ becomes an ISSS set. Furthermore, for any compact set $E$, the minimal condensation set $C=\overline{E\setminus\bigcup_{i=1}^{N}f_{i}(E)}$ ensures that $E$ is an ISSS set.
@@ 4. Construction of Inhomogeneous Sub-Self-Similar Sets
$O_{S}:=\bigcup_{{\omega}\in S^{*}}f_{{\omega}}(C).$
@@ 5.1. Hausdorff dimension
Let $E\cup O_{S}$ be an ISSS set with SSS set $E$ and condensation set $C.$ Then $\dim_{H}(E\cup O_{S})=\max\{s,\dim_{H}C\}$.
@@ 5.2. Upper box dimension
The function $\tau(s)=\lim_{k\to\infty}\bigg(\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\bigg)^{1/k}$ defined in Theorem 5.1 is strictly decreasing continuous function with $0\leq\tau(s)<\infty$ and $\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\geq\tau(s)^{k}$ for all $k=1,2,\ldots$
@@
By Theorem 5.1 $\overline{\dim}_{B}E\leq s,$ it is sufficient to show that $\overline{\dim}_{B}O_{S}\leq\max\{s,\overline{\dim}_{B}C\}.$ Let $\max\{s,\overline{\dim}_{B}C\}<t,$ then by the definition of the box-dimension, there exists some constant $c_{t}$ such that
@@
The third line comes from Lemma 5.6. The sixth line follows because $X$ is a compact space and $\delta/\rho_{\omega}>1$, as established by Lemma 5.4. This completes the proof. ∎
@@ 5.3. Lower box dimension
By Lemma 5.4, for $t>s$, the cardinality of the $\delta$-stopping set $S$ is given by:
@@
Let $t>\max\{s,\underline{\dim}_{B}C\}$. Then, by Lemma 5.8, we have $P_{t}(C)\leq\underline{\dim}_{B}C<t$. Thus, by Lemma 5.9, for all $\epsilon\in(0,1-P_{t}(C))$, there exists $\delta\in(0,\epsilon)$ such that:
@@ 6. Continuity of the Hausdorff dimension of ISSS
Let $S^{k}:=\{\omega=\omega_{1}\omega_{2}\ldots\omega_{k}:\exists\omega_{1}\ldots\omega_{n}\ldots\in S\}.$ Define IFSs $I_{k}:=\{f_{\omega}:\omega\in S^{k}\}$ with corresponding ISSS set $F_{k}$ for each $k\in\mathbb{N}$. Then, For any $x\in F=E\cup O_{S}$, we consider two cases:
@@
For each $k\in\mathbb{N}$, we rewrite $\omega$ as $\omega:=(\omega_{1}\omega_{2}\ldots\omega_{k})|_{0}(\omega_{k+1}\omega_{k+2}\ldots\omega_{2k})|_{1}\ldots(\omega_{jk+1}\omega_{jk+2}\ldots\\ \omega_{(j+1)k})|_{j}\ldots\in S.$ Since, $S$ is closed under left shift, it implies that each block $\omega_{j}:=\omega_{jk+1}\ldots\omega_{(j+1)k}\in S^{k}$ for each $j=0,1,\ldots$, and we have
$x=\lim_{j\to\infty}f_{\omega_{0}}\circ f_{\omega_{1}}\circ\ldots\circ f_{\omega_{j}}(z),$
@@
Conversely, suppose $x\in\bigcup_{k=1}^{\infty}F_{k}\cup O_{S}$. If $x\in O_{S}$, then the claim holds trivially. Now suppose $x\in F_{k}$ for some $k$. Then there exists a sequence $\{\omega_{j}\}_{j=1}^{\infty}$ in $S^{k}$ such that
$x=\lim_{j\to\infty}f_{\omega_{1}}\circ f_{\omega_{2}}\circ\ldots\circ f_{\omega_{j}}(z).$
Using the fact that $\Theta(S)=E$ and that the concatenation $\omega=\omega_{1}\omega_{2}\ldots\omega_{j}\ldots\in S$, we conclude that $x\in F$.
@@
Note 6.2.
It is worth noting that, if the underlying IFS satisfies the OSC, then the IFS $I_{k}$ also satisfy the OSC for each $k\in\mathbb{N},$ and $\{E_{k}\}_{{k=1}^{\infty}}$ is a monotonic increasing sequence. If $x\in E_{k},$ then there exists a sequence $\{\omega_{j}\}_{j=1}^{\infty}$ in $S^{k}$ such that $\lim_{j\to\infty}f_{\omega_{1}}\circ f_{\omega_{2}}\circ\cdots\circ f_{\omega_{j}}(z)=x.$ By the fact that $x\in F$ and $\Theta(S)=E,$ the concatenation $\Omega:=\omega_{1}\ldots\omega_{j}\ldots\in S.$ Since, $\omega_{j}\in S^{k},$ by our construction there exist $\omega^{j}\in S$ such that $\omega^{j}|_{k}=\omega_{j}:=\omega^{j}_{1}\ldots\omega^{j}_{k}$ for each $j\in\mathbb{N}.$ Now, we rewrite $\Omega=(\omega^{1}_{1}\ldots\omega^{1}_{k}\omega^{2}_{1})(\omega^{2}_{2}\ldots\omega^{2}_{k}\omega^{3}_{1}\omega^{3}_{2})\ldots$, then clearly each block of length $k+1$ and $S^{k+1}.$ This implies that $x\in E_{k+1}.$
@@
Let $F=E\cup O_{S}$ be the ISSS set corresponding to the IFS $\{f_{i}\}_{i=1}^{N}.$ Then, there exists a sequence of IFSs $\{I_{k}\}_{k=1}^{\infty}$ with ISSS sets $\{F_{k}\}$ and under inhomogeneous open set condition (IOSC) $\dim_{H}E=\lim_{k\to\infty}\dim_{H}E_{k}=\lim_{k\to\infty}s_{k},$ where $s_{k}$ is the Hausdorff dimension of $F_{k}$ defined in Lemma 6.3.
@@
By Theorem 6.1, there exists a sequence of IFSs $\{I_{k}\}_{k=1}^{\infty}$ with ISSS sets $F_{k}$ such that $F=\bigcup_{k=1}^{\infty}F_{k}.$ It is easy to observe that if $\{f_{i}\}_{i=1}^{N}$ satisfies the IOSC, then each $I_{k}$ also satisfies the IOSC. If $s_{k}$ represents the Hausdorff dimension of $E_{k}$, which satisfies